Sol Summary – February 1st

Crew 188 Sol 4 Summary Report 1FEB2018

Sol 4

Summary Title: A sol featuring soul with the devastation of soles Author’s name: Ryan L. Kobrick, Ph.D., MDRS Crew 188 Commander Mission Status: All nominal

Sol Activity Summary:
Let’s jump to the bottom line. My hiking boots were completely destroyed on our hike to Phobos Peak with the soles peeling off the bottom. An EVA with modest to difficult objectives was challenged with equipment failures and ingenuity to solve problems as they occurred. This is the type of exploration we need for a Mars mission. Calculating risks, making decisions, and with collective exploration experience solving problems. There’s going to be a lot of problems. The EVA Macgyver of the day was Sarah Jane, fixing one boot with a safety whistle, and the other with my laces and some zip ties I was using. Between Julia, SJ and I we were challenged physically and mentally. I had suggested Phobos Peak for the EVA, and I was clearly willing to lose my boots over it. Luckily my hiking shoes are rugged and ready to step up to the next EVA challenge.

The crew enjoyed the scenic views from the science dome during a session of Kung Fu Thai Chi in the morning. Every morning needs to start off on the right foot… even if part of your foot falls off later. The crew also cracked into Cottage Cheese with Blackcurrant Puree space desert from the Russian Space Food Laboratory. It’s apparently “approved by space”. Further work continued on the NASA VEGGIE OASYS project in the Green Hab. We also finally opened the dust load experiment that I am running from NASA Glenn Research Center. I’m excited to set up the hardware and see how much dust we are tracking into the hab. It’s a dusty environment, and knowing how dirty it gets during key activities (EVA airlock, post EVA clean up, and more) we can help prepare future life support systems with the correct type of filters.

Spacesuit Up!
Ryan L. Kobrick, Ph.D.
MDRS Crew 188 Commander

Look Ahead Plan:
EVA 5 is planned for tomorrow to revisit a site from EVA 2 and to further add navigation and waypoint data for the crew to use on all EVAs. Meanwhile several other projects will be underway inside the dome.

Anomalies in work:
MDRS Director is coordinating power system updates with contractors. There was a site visit today.
Robotic observatory currently not functional. MDRS Astronomy lead working problem.

The weather was clear until mid-day when clouds rolled into the area.

Crew Physical Status:
It was a busy day all around, so we are looking at our mission ahead to plan out key EVAs, science projects, and fun activities to make sure we are having a well-balanced experience.

One is on the books for tomorrow.

Reports to file:
1. EVA #5 requests
2. Ops Report
3. Sol Summary
4. Journalist Report
5. Creative Report
6. Green Hab Report
7. Daily Photos

Support Requested:
Website updates.
Microwave timer works, but main chamber is not functioning/heating. [end]

Copyright © The Mars Society. All rights reserved. | Main Site